OpenStudy (nani200):

find the domain of the given function f(x) = sqrt x + 3 / (x+ 8)(x-2) A. x > 0 B. all real numbers C. x>/= -3, x cannot = 2 D. x cannot = 8, x cannot = -3, x cannot = 2

myininaya (myininaya):

\[f(x)=\sqrt{x}+\frac{3}{(x+8)(x-2)} ?\]

OpenStudy (nani200):

the sqrtx is in the denominator with the 3

myininaya (myininaya):

it's in the denominator ?

myininaya (myininaya):

i'm not sure how to interpret it then the denominator is in bottom

OpenStudy (nani200):

sorry i meant numerator

OpenStudy (nani200):

the numerator is sqrt x + 3

myininaya (myininaya):

\[f(x)=\frac{\sqrt{x}+3}{(x+8)(x-2)}\]

OpenStudy (nani200):

yes

myininaya (myininaya):

first do you know the domain of the function g(x)=sqrt(x) ?

OpenStudy (nani200):

no

myininaya (myininaya):

what happens when you have sqrt(negative number)?

OpenStudy (nani200):

i don't think you can do that

myininaya (myininaya):

so you can only do square root of positive or zero numbers (well over the real numbers anyways)

myininaya (myininaya):

so the domain for g(x)=sqrt(x) is x is positive or zero we say x is in [0,inf) where inf means infinity

myininaya (myininaya):

so now let's look at your denominator

myininaya (myininaya):

we cannot divided by zero what values of x will make the denominator 0?

OpenStudy (nani200):

-8 and +2

myininaya (myininaya):

right so the domain of the function \[f(x)=\frac{\sqrt{x}+3}{(x+8)(x-2)} \text{ is } \\ x \ge 0 \text{ but } x \neq -8 \text{ or } 2 \\ \text{ we don't even have \to mention } x \neq -8 \text{ since that is redundant } \\ \text{ so you can say instead } \\ x \ge 0 \text{ but } x \neq 2\]

myininaya (myininaya):

are you sure the numerator wasn't suppose to be sqrt(x+3) ?

OpenStudy (nani200):

the sqrt sign is over the x and the 3

myininaya (myininaya):

so it was suppose to be sqrt(x+3) and not sqrt(x)+3 \[f(x)=\frac{\sqrt{x+3}}{(x+8)(x-2)}\]

OpenStudy (nani200):

right

myininaya (myininaya):

use what we said about square roots above you need the inside to be positive or zero of the square root so you need x+3 to need be positive or zero in math symbols that looks like \[x+3 \ge 0\] solve this and then you already pointed out that x cannot be -8 or 2 so anyways let me know what you get when you solve the above inequality

OpenStudy (nani200):

x is greater than or equal to -3

myininaya (myininaya):

right now from that set we want to exclude -8 and 2 one of these values is already excluded from the set x> or =-3

OpenStudy (nani200):

so its d

OpenStudy (nani200):

?